MICROWAVE SPECTRA OF THE Xe-N_{2} VAN DER WAALS COMPLEX: A COMPARISON OF EXPERI-MENT AND THEORY

\underline{QING WEN} and WOLFGANG JÄGER, Department of Chemistry, University of Alberta, Edmonton, AB T6G 2G2, Canada.

Rotational transitions of the Xe-N_{2} complex were measured in the 4-18 GHz frequency region using a pulsed-nozzle Fourier-transform microwave spectrometer. Twelve (four) a-type transitions were recorded for the ^{132}Xe-^{14}N_{2} and ^{129}Xe-14N_{2} (^{131}Xe-^{15}N_{2}) isotopomers. In addition, the nuclear quadrupole hyperfine structures due to the presence of the ^{14}N (I=1) and ^{131}Xe (I=3/2) nuclei were detected and analyzed. A high level ab initio potential energy surface was calculated at the CCSD(T) level of theory. Well-Tempered Basis Set (WTBS) with additional polarization functions was used for the Xe atom and aug-cc-pVTZ basis set for the N atoms. The basis sets were supplemented with midbond functions. This surface has a global minimum at a T-shaped geometry with a well depth of -122.6 cm^{−1}. Bound state energies supported by the potential energy surface were determined using the JACOBI computer code of X. Song and P. N. Roy. The quality of the ab initio potential energy surface is evaluated by comparison of the experimental transition frequencies and rotational and centrifugal distortion constants with those derived from the bound state energies. A scaled potential energy surface was obtained which has excellent agreement with the experimental data.